# LLLMs: A Data-Driven Survey of Evolving Research on Limitations of Large Language Models

### v2

@@ 4.4.2. Trend Analysis
ii arXiv Topic Distribution Per Year (LlooM clustering approach)
We begin by examining how visible different limitation topics are across the broader LLM research field. Figure 9 shows the annual distribution of LLM limitation topics across ACL and arXiv, normalized by all LLM-focused papers. These proportions reflect the overall visibility of each topic. Additionally, to capture how visibility changes over time, we compute the relative percentage change in LLM-normalized topic share from year $y$ to $y+1$, defined as $\frac{\text{Share}_{y+1}-\text{Share}_{y}}{\text{Share}_{y}}\times 100$, where $\text{Share}_{y}$ refers to the LLM-wide share of a given topic in year $y$, i.e., the proportion of all LLM papers that address that topic (see Table 16 in supplementary material).
@@
  • $\downarrow$ Decreasing: No topics show statistically significant declines. However, Bias and Fairness drops from $\sim$17to 10%, and Language and Cultural Limitations from $\sim$20to 12% between 2023-Q2 and 2024-Q4. Computational Cost, after a steady rise in the period from late 2022 to late 2023, decreases from $\sim$5in the late 2023 to below 5% by 2024-Q4. None of these changes are statistically significant.
@@
  • $\rightarrow$ Stable or Fluctuating: Most remaining topics show no significant directional movement. Reasoning stays around 20–30%, Trustworthiness holds at 40–45%, and Overconfidence, Long Context, and Privacy Risks remain within narrow ranges. Prompt Sensitivity declines from 10% to $\sim$5(2023–2025), while Hallucination and Language and Cultural Limitations rise modestly and then level off. Catastrophic Forgetting stays consistently below 4%.
@@
The stabilization of topic trends coincides with a sharp rise in raw paper counts beginning in 2023 (see Figure 16), indicating not just increased research volume, but a shift toward a more coherent field. Before 2023-Q2, most topics appear in fewer than 25 ACL papers and under 100 in arXiv, making early signals harder to interpret. This growth aligns with the release of ChatGPT in November 2022, as well as the emergence of other major models like GPT-4 (Achiam et al., 2023), PaLM (Chowdhery et al., 2023), and LLaMA (Touvron et al., 2023) between February and July 2023, which likely contributed to the expansion and differentiation of LLM limitations research during this period.
@@ 4.4.3. LLLMs Topics Distribution Across ArXiv Categories
Figure 12. Limitation topic trends across four main arXiv categories (cs.CL, cs.LG, cs.AI, cs.CV). To maintain visual clarity, only the five most frequent topics (based on their overall percentage distribution across categories, as shown in Figure 18 in the supplementary material) are shown; others are grouped as ?Other.? Grey shading in cs.AI and cs.CV marks periods before 2023 with insufficient data for these categories.
Although many categories share topics such as Trustworthiness, Reasoning, Generalization, and Alignment Limitations, both their overall topic composition and temporal dynamics vary by field. Figure 18 (right) in the supplementary material, shows how topic shares differ across arXiv categories, while Figure 12 illustrates how topics in the four largest categories evolve over time. Key trends in the largest categories are as follows:
@@
Figure 13. Distribution of limitations-related topics in six arXiv categories with lower paper counts in our dataset, as shown in Figure 18. Each chart includes only papers where the category is assigned as the primary arXiv category. Topics that make up less than 3% are grouped under Other.
@@
4.5. Method and Output Comparison: HDBSCAN vs. LlooM
The aforementioned topic distributions differ depending on the clustering method. In this section, we compare the two approaches used, HDBSCAN+BERTopic (see Section C in the supplementary material) and LlooM (see 4.4), to identify which findings are stable and which may be method-specific. We first compare trend agreement, then examine methodological differences to better explain any observed divergences in their results.
To compare the trends, we report (i) Kendall’s Tau for each individual topic (as determined in the trend analyses in Section C in the supplementary material and 4.4) to assess alignment in trend direction and significance, and (ii) the Spearman correlation of the time series between matched topics from HDBSCAN and LlooM to assess the similarity of overall trend shapes. To identify matching topics between HDBSCAN and LlooM, we select the best match for each cluster based on identical or semantically similar names. To validate these matches, we compute the Jaccard overlap between their associated paper sets. Specifically, for each cluster produced by HDBSCAN, we compute its Jaccard similarity with all LlooM topics, and vice versa. Given two paper sets $X$ from HDBSCAN and $Y$ from LlooM, the Jaccard similarity is defined as $\text{Jaccard}(X,Y)=\frac{|X\cap Y|}{|X\cup Y|}$. We select the top-1 Jaccard score for each cluster, representing its highest similarity with any topic from the other method. Best-matching topic pairs are shown in Table 17 and Table 18 in the supplementary material, confirming that major topics identified by name also show substantial paper overlap.
Trend Alignment Between HDBSCAN and LlooM
Table 10. Comparison of limitation trends identified by HDBSCAN and LlooM for ACL and arXiv datasets. Trend direction ( increasing, flat, decreasing) is based on Kendall’s Tau from the Mann-Kendall test. Significance is indicated with an asterisk (*) and reported only for increasing or decreasing trends. Match symbols: $=$ full agreement (direction and significance), $=$ partial agreement (direction only), ×$=$ disagreement.
@@
[…]
@@
These results suggest that while LlooM and HDBSCAN identify similar broad limitation areas, they organize papers differently at a finer-grained level. This pattern is further supported by Table 17 and Table 18 in the supplementary material, which show the most closely aligned topics across methods. Some large limitation areas appear relatively stable across clustering strategies: e.g. topics such as Reasoning, Hallucination, Security Risks, and Bias and Fairness align well across datasets and clustering approaches ($J>0.4$), whereas smaller topics, such as Overconfidence and Prompt Sensitivity, are often merged into broader categories (e.g., Hallucination). This reflects differences in clustering granularity: LlooM tends to split topics into overlapping subcategories, while HDBSCAN merges related issues into broader clusters. For instance, HDBSCAN combines LlooM’s Security Risks, Privacy Risks, and Trustworthiness into a single Security topic.
@@ 2. Within LLLMs research, reasoning is the most frequent topic, but research is diverse
Beyond these, LLM limitation research is notably diverse. Our clustering analyses (HDBSCAN and LlooM) reveal a broad spectrum of concerns, ranging from code generation and benchmark contamination to prompt sensitivity and long context. This breadth reflects the current state of LLM limitation research: a fast-growing, methodologically diverse field still defining its major challenges. Additionally, as shown in Section D in the supplementary materials, many papers address multiple limitations simultaneously, reflecting the complexity of emerging concerns.
@@
These trends align with shifting priorities in the LLM community. The growing attention to alignment and security reflects their increasingly central role in both training and evaluation of LLMs. Though still relatively new and unsettled (Shen et al., 2023), these concerns became prominent in 2022 with the rise of Reinforcement Learning from Human Feedback (RLHF), which is now foundational in the training pipelines of major models (Wang et al., 2024a). Yet ensuring safety without compromising performance remains an open challenge (Qi et al., 2023; Tan et al., 2025), making this an active and fast-moving research area.
3.2. Limitations persist as applications expand. As LLMs are deployed in high-stakes domains, interest in hallucination and knowledge editing is growing due to the increasing demand for factual accuracy and controllability. These remain deeply challenging: hallucination is increasingly seen as an inherent property of LLMs, rooted in model architecture itself (Banerjee et al., 2024; Xu et al., 2024a).
Finally, these challenges grow as models move beyond text. The rise of multimodality-related limitations suggests that LLMs not only inherit existing issues but also encounter new ones with inputs like images and audio (Zhang et al., 2024). This trend likely reflects growing interest caused by the release of GPT-4V (Achiam et al., 2023), LLaVA (Liu et al., 2023), and other vision-language models in mid-2023. This and other trends discussed above coincide with broader shifts already noted in previous studies of the arXiv corpus: from early 2023 to late 2024, top-cited LLM papers increasingly came from cs.CV and cs.LG, with cs.CL seeing a relative decline (Leiter et al., 2024). Similarly, authorship diversified, with many newcomers from computer vision, security, and software engineering (Movva et al., 2023).
4. Despite methodological differences, HDBSCAN and LlooM identify overlapping high-frequency topics (e.g., Reasoning, Hallucination, Security Risks) and similar trend patterns, pointing to the stability of the main findings.
We validate our results by comparing HDBSCAN+BERTopic (single-topic, density-based) with LlooM (LLM-based, multi-topic). Despite methodological differences, both clustering approaches identify the high-frequency topics, most notably Reasoning, Hallucination, and Security Risks, showing substantial agreement in both topic composition and trend trajectories. And although smaller topics (e.g., Prompt Sensitivity) and trend significance can vary, the main trends reported in this study appear stable across methods.
@@
  • Both of our clustering approaches are prone to some instability. LlooM can be variable due to its reliance on LLM outputs, a limitation noted by its authors as well (Lam et al., 2024). HDBSCAN+BERTopic might also show some run-to-run variability, stemming from the stochastic nature of UMAP and sensitivity to embedding changes. While high-level patterns are generally stable, topic composition and temporal trends may shift slightly. We mitigate these issues by validating results across both methods.
  • While we adopt a broad definition of LLMs in both automated and human annotation (including transformer-based, foundational, multimodal models), this scope may still introduce bias, e.g., we may not capture newly emerging terms in fields such as computer vision.
@@
Future work could refine and extend these findings in several directions. First, the limitation topics could be decomposed into subcategories, such as types of reasoning or specific forms of bias, using hierarchical or agglomerative clustering techniques. Second, extending the analysis to earlier years, especially the period following BERT’s introduction in 2018, could clarify how concerns raised for smaller-scale PLMs have shifted, declined, or re-emerged with model scaling.
@@ Appendix A Keyword List
Table 13. Representative examples from the human rating annotation task, with brief explanation for abstracts’ ratings on the 0–5 scale (defined in Table 2). In higher-rated papers (3–5), evidence of LLLMs is highlighted in bold.
@@
3 ?[…] it remains difficult to distinguish effects of statistical correlation from more systematic logical reasoning grounded on the understanding of real world [in PLMs]. […] We find that models are consistently able to override real-world knowledge in counterfactual scenarios… however, we also find that for most models this effect appears largely to be driven by simple lexical cues. […]? (Li et al., 2023d), ACL 2023 Discusses reasoning limitations in moderate detail, but not as the main focus.
@@
Figure 14. Prompt 3 used for LLM-based classification and evidence extraction (as discussed in 4.4.3). Abstract texts and evidence are shortened for brevity).

### v3

@@ 4.3.2. Trend Analysis
$i$ arXiv Topic Distribution Per Year (LlooM clustering approach)
We begin by examining how visible different limitation topics are across the broader LLM research field. Figure 8 shows the annual distribution of LLM limitation topics across ACL and arXiv, normalized by all LLM-focused papers. These proportions reflect the overall visibility of each topic. Additionally, to capture how visibility changes over time, we compute the relative percentage change in LLM-normalized topic share from year $y$ to $y+1$, defined as $\frac{\text{Share}_{y+1}-\text{Share}_{y}}{\text{Share}_{y}}\times 100$, where $\text{Share}_{y}$ refers to the LLM-wide share of a given topic in year $y$, i.e., the proportion of all LLM papers that address that topic (see Table 5 in the online supplementary material).
@@
  • $\downarrow$ Decreasing: No topics show statistically significant declines. However, Bias and Fairness drops from $\sim$17% to 10%, and Language and Cultural Limitations from $\sim$20% to 12% between 2023-Q2 and 2024-Q4. Computational Cost, after a steady rise in the period from late 2022 to late 2023, decreases from $\sim$5% in the late 2023 to below 5% by 2024-Q4. None of these changes are statistically significant.
@@
  • $\rightarrow$ Stable or Fluctuating: Most remaining topics show no significant directional movement. Reasoning stays around 20–30%, Trustworthiness holds at 40–45%, and Overconfidence, Long Context, and Privacy Risks remain within narrow ranges. Prompt Sensitivity declines from 10% to $\sim$5% (2023–2025), while Hallucination and Language and Cultural Limitations rise modestly and then level off. Catastrophic Forgetting stays consistently below 4%.
@@
The stabilization of topic trends coincides with a sharp rise in raw paper counts beginning in 2023 (see Figure 3 in the online supplementary material), indicating not just increased research volume, but a shift toward a more coherent field. Before 2023-Q2, most topics appear in fewer than 25 ACL papers and under 100 in arXiv, making early signals harder to interpret. This growth aligns with the release of ChatGPT in November 2022, as well as the emergence of other major models like GPT-4 (Achiam et al., 2023), PaLM (Chowdhery et al., 2023), and LLaMA (Touvron et al., 2023) between February and July 2023, which likely contributed to the expansion and differentiation of LLM limitations research during this period.
@@ 4.3.3. LLLMs Topics Distribution Across ArXiv Categories
Figure 11. Limitation topic trends across four main arXiv categories (cs.CL, cs.LG, cs.AI, cs.CV). To maintain visual clarity, only the five most frequent topics (based on their overall percentage distribution across categories, as shown in Figure 5 in the online supplementary material) are shown; others are grouped as ?Other.? Grey shading in cs.AI and cs.CV marks periods before 2023 with insufficient data for these categories.
Figure 12. Distribution of limitations-related topics in six arXiv categories with lower paper counts in our dataset, as shown in Figure 5 in the online supplementary material. Each chart includes only papers where the category is assigned as the primary arXiv category. Topics that make up less than 3% are grouped under Other.
Although many categories share topics such as Trustworthiness, Reasoning, Generalization, and Alignment Limitations, both their overall topic composition and temporal dynamics vary by field. Figure 5 (right) in the online supplementary material, shows how topic shares differ across arXiv categories, while Figure 11 illustrates how topics in the four largest categories evolve over time. Key trends in the largest categories are as follows:
@@
4.4. Clustering Methods Comparison
The aforementioned topic distributions differ depending on the clustering method. In this section, we compare HDBSCAN+BERTopic (Section C in the online supplementary material) and LlooM (4.3.2) to identify which findings are stable and which may be method-specific. We first assess trend agreement, then explore methodological differences to explain any result divergences.
4.4.1. Cross-Method Comparison Between HDBSCAN and LlooM
Table 8. Comparison of limitation trends identified by HDBSCAN and LlooM for ACL and arXiv datasets. Trend direction ( increasing, flat, decreasing) is based on Kendall’s Tau from the Mann-Kendall test. Significance is indicated with an asterisk (*) and reported only for increasing or decreasing trends. Match symbols: == full agreement (direction and significance), == partial agreement (direction only), ×== disagreement.
@@ 4.4.1. Cross-Method Comparison Between HDBSCAN and LlooM
$\dagger$} Topics are matched by Jaccard overlap, except for Knowledge Editing, which was manually aligned based on topic names due to weak overlap.
To compare the trends, we report (i) Kendall’s Tau for each individual topic (as determined in the trend analyses in Section C in the online supplementary material and 4.3.2) to assess alignment in trend direction and significance, and (ii) the Spearman correlation of the time series between matched topics from HDBSCAN and LlooM to assess the similarity of overall trend shapes. To identify matching topics between HDBSCAN and LlooM, we first identify candidate matches for each cluster based on identical or semantically similar names and validate these by computing the Jaccard overlap between their associated paper sets [^6] All matches except Knowledge Editing were confirmed by highest Jaccard overlap; Knowledge Editing was aligned manually. Specifically, for each cluster produced by HDBSCAN, we compute its Jaccard similarity with all LlooM topics, and vice versa. Given two paper sets $X$ from HDBSCAN and $Y$ from LlooM, the Jaccard similarity is defined as $\text{Jaccard}(X,Y)=\frac{|X\cap Y|}{|X\cup Y|}$. We select the top-1 Jaccard score for each cluster, representing its highest similarity with any topic from the other method. Best-matching topic pairs are shown in Tables 9 and 10 in the online supplementary material, confirming that major topics identified by name also show substantial paper overlap.
Trend Alignment Between HDBSCAN and LlooM
@@
These results show that LlooM and HDBSCAN identify similar broad limitation areas but organize papers differently at a finer-grained level. This is supported by Tables 9 and 10 in the online supplementary material, which report the closest topic alignments between methods. Large areas such as Reasoning, Hallucination, Security Risks, and Bias and Fairness appear relatively stable and align well across methods and datasets ($J>0.4$), whereas smaller topics, including Overconfidence and Prompt Sensitivity, are often absorbed into broader categories such as Hallucination. This reflects differences in clustering granularity: LlooM splits topics into overlapping subcategories, while HDBSCAN merges related issues into broader clusters. For example, HDBSCAN merges LlooM’s Security Risks, Privacy Risks, and Trustworthiness into a single Security cluster.
@@
4.4.2. Human Evaluation
While a full gold-standard clustering is not scalable, we assess cluster validity through stratified human re-annotation of cluster assignments and measure both inter-annotator and human–method agreement for both HDBSCAN and LlooM.
For both ACL and arXiv, we construct a stratified human evaluation dataset (50 papers for ACL and 60 for arXiv) by sampling papers with approximately uniform per-topic sample sizes, including outliers, stratified with respect to HDBSCAN clusters. Each sampled paper is annotated with both its HDBSCAN and LlooM topic assignments. Because LlooM allows multi-label assignments, LlooM topic frequencies in the evaluation dataset are not explicitly controlled during sampling; however, all LlooM topics identified in the corpus are represented. [^8] Except for the broad Trustworthiness cluster in arXiv, which was excluded from human annotation due to substantial overlap with other topics.
The dataset was annotated by a professor (machine learning) and a PhD student (NLP). Annotators were provided with each paper’s title, extracted evidence, and keyphrases, corresponding to the input used by the clustering methods (as described in Section 3.5). For HDBSCAN, annotators selected a single topic per paper. For LlooM, annotators could assign multiple topics per paper, reflecting its multi-label clustering setup. Annotators were shown representative BERTopic keyword lists and LlooM assignment prompts to convey cluster semantics.
We measure both inter-annotator agreement and human–method agreement, using Cohen’s $\kappa$ for HDBSCAN, which produces single-label assignments, and boot-F1 (Marchal et al., 2022) for LlooM to account for chance agreement in the multi-label setting.
Table 10. Inter-annotator and human–method agreement for HDBSCAN and LlooM on ACL and arXiv datasets. Cohen’s $\kappa$ is used for HDBSCAN (single-label), and boot-F1 is used for LlooM (multi-label).
Dataset Method Inter-Annotator Human–Method
ACL HDBSCAN $\kappa=0.79$ $\kappa=0.74$
ACL LlooM boot-F1 =0.70=0.70 boot-F1 =0.61=0.61
arXiv HDBSCAN $\kappa=0.71$ $\kappa=0.55$$0.59$
arXiv LlooM boot-F1 =0.53=0.53 boot-F1 =0.49=0.490.510.51
Table 10 shows substantial inter-annotator agreement for both datasets ($\kappa=0.71$$0.79$ for HDBSCAN and boot-F1 =0.53=0.530.700.70 for LlooM), indicating consistent human judgments across both single-label and multi-label settings. Human–method agreement is also substantial for ACL ($\kappa=0.74$, boot-F1 =0.61=0.61) and moderate for arXiv ($\kappa=0.55$$0.59$, boot-F1 =0.49=0.490.510.51), indicating that human annotations broadly correspond to the topics produced by the clustering methods.
@@ 2. Within LLLMs research, reasoning is the most frequent topic, but research is diverse
Beyond these, LLM limitation research is notably diverse. Our clustering analyses (HDBSCAN and LlooM) reveal a broad spectrum of concerns, ranging from code generation and benchmark contamination to prompt sensitivity and long context. This breadth reflects the current state of LLM limitation research: a fast-growing, methodologically diverse field still defining its major challenges. Additionally, as shown in Section D in the online supplementary material, many papers address multiple limitations simultaneously, reflecting the complexity of emerging concerns.
@@
These trends align with shifting priorities in the LLM community. The growing attention to alignment and security reflects their increasingly central role in both training and evaluation of LLMs. Though still relatively new and unsettled (Shen et al., 2023), these concerns became prominent in 2022 with the rise of Reinforcement Learning from Human Feedback (RLHF), which is now foundational in the training pipelines of major models (Wang et al., 2024d). Yet ensuring safety without compromising performance remains an open challenge (Qi et al., 2023; Tan et al., 2025), making this an active and fast-moving research area.
3.2. Limitations persist as applications expand. As LLMs are deployed in high-stakes domains, interest in hallucination and knowledge editing is growing due to the increasing demand for factual accuracy and controllability. These remain deeply challenging: hallucination is increasingly seen as an inherent property of LLMs, rooted in model architecture itself (Banerjee et al., 2024; Xu et al., 2024).
Finally, these challenges grow as models move beyond text. The rise of multimodality-related limitations suggests that LLMs not only inherit existing issues but also encounter new ones with inputs like images and audio (Zhang et al., 2024). This trend likely reflects growing interest caused by the release of GPT-4V (Achiam et al., 2023), LLaVA (Liu et al., 2023), and other vision-language models in mid-2023. These and other trends discussed above coincide with broader shifts already noted in previous studies of the arXiv corpus: in particular, previous work finds that from early 2023 to late 2024, top-cited LLM papers increasingly came from cs.CV and cs.LG, with cs.CL seeing a relative decline (Leiter et al., 2024). Similarly, authorship diversified, with many newcomers from computer vision, security, and software engineering (Movva et al., 2023).
4. Despite methodological differences, HDBSCAN and LlooM identify overlapping high-frequency topics (e.g., Reasoning, Hallucination, Security Risks) and similar trend patterns, supporting the stability of the main findings.
We validate our results by comparing HDBSCAN+BERTopic (single-topic, density-based) with LlooM (LLM-based, multi-topic). Despite their differences, both methods recover the same dominant topics, especially Reasoning, Hallucination, and Security Risks, with strong agreement in topic composition and trend trajectories. And although smaller topics (e.g., Prompt Sensitivity) and trend significance can vary, the main trends appear stable across methods.
@@
  • Both of our clustering approaches are prone to some instability. LlooM can be variable due to its reliance on LLM outputs, a limitation noted by its authors as well (Lam et al., 2024). HDBSCAN+BERTopic can also vary across runs because UMAP is stochastic and sensitive to embedding changes. While high-level patterns are generally stable, topic composition and temporal trends may shift slightly. We mitigate these issues by validating results across both methods.
  • While we adopt a broad definition of LLMs in both automated and human annotation (including transformer-based, foundational, multimodal models), this scope may still introduce bias. In particular, our reliance on abstract- and keyword-level coverage and on the extraction of explicitly stated limitation evidence may undercount limitations discussed using non-standard terminology, or newly emerging terms (e.g., in fields such as CV).
@@
  • Given our automated survey design, trend analyses reflect the prevalence of limitation-related discussions in research rather than the severity of the discussed limitations. Although sustained research interest can reflect practical relevance, changes in topic prevalence should primarily be interpreted as shifts in research attention.
  • This survey provides a large-scale overview of LLLMs research rather than a fine-grained taxonomy; while the identified categories do not capture detailed limitation subtypes, the topics are high-level to reflect dominant and stable themes across methods and datasets.
Future work can extend these findings in several ways. First, limitation topics can be decomposed into finer subcategories, such as specific reasoning types or bias forms, using hierarchical or agglomerative clustering. Second, expanding the analysis to earlier years, especially after BERT’s introduction in 2018, could show how concerns about smaller PLMs evolved with model scaling. Third, the automatic pipeline can be adapted to incorporate newly published papers, enabling regular updates of the dataset over time.
Acknowledgements.
This research was funded by the Ministry of Culture and Science of the State of North Rhine-Westphalia under the grant no NW21-059A (SAIL). This project originated at the 2024 retreat of the Natural Language Learning & Generation (NLLG) Lab at TU Nürnberg. We thank Yanran Chen, Christoph Leiter, Ran Zhang, Daniil Larionov, and Jonas Belouadi for valuable early input and discussions that helped shape this work.
@@ Appendix A Keyword List
Table 12. Representative examples from the human rating annotation task, with brief explanation for abstracts’ ratings on the 0–5 scale (defined in Table 2 in the main paper). In higher-rated papers (3–5), evidence of LLLMs is highlighted in bold.
@@
3 ?[…] it remains difficult to distinguish effects of statistical correlation from more systematic logical reasoning grounded on the understanding of real world [in PLMs]. […] We find that models are consistently able to override real-world knowledge in counterfactual scenarios… however, we also find that for most models this effect appears largely to be driven by simple lexical cues. […]? (Li et al., 2023a), ACL 2023 Discusses reasoning limitations in moderate detail, but not as the main focus.
@@
Figure 13. Prompt 3 used for LLM-based classification and evidence extraction (as discussed in Section 3.4 of the main paper). Abstract texts and evidence are shortened for brevity.